OpenStudy (anonymous):

A rhombus and a rectangle have equal perimeters. The rhombus has side 15 cm. The length of the rectangle is 6 cm more than the width. What is the are if the rectangle?

OpenStudy (anonymous):

perimetre of rhombus=60 therfore perimetre of rectangle =60 Rectangle: p=2l +2(l+6) 60=2l+2l +12 48=4l l=12 therfore (12 +6)=18 width =18 A= 12 x 18 =216 Hope this helps. If you need it explained any more just ask:)